OpenStudy (anonymous):

The equation y = 350x + 480 describes the amount of money in Blake’s savings account, where y is the amount of money in his account and x is the number of monthly deposits he makes. Eddie also has a savings account, and the balances in his account are recorded in the table. How much money did each account start with? Number of Monthly Deposits 1 2 4 7 Balance ($) 900 1375 2325 3750 Blake’s account started with $ . Eddie’s account started with $ .

OpenStudy (anonymous):

I'm Stuck

OpenStudy (rocketisawesome):

ok I will help you

OpenStudy (anonymous):

ty

OpenStudy (vickyisthesmartone123):

please medal me i need one more to get to 50

OpenStudy (rocketisawesome):

hmmm

OpenStudy (vickyisthesmartone123):

pwease?

OpenStudy (rocketisawesome):

Fine i will medal you if you help me with one question

OpenStudy (vickyisthesmartone123):

ok i will

OpenStudy (vickyisthesmartone123):

hurry please! :)

OpenStudy (rocketisawesome):

there is ur medal

OpenStudy (anonymous):

@Vickyisthesmartone123 can you help me I really need help

OpenStudy (vickyisthesmartone123):

Ok let me see here...

OpenStudy (anonymous):

ok

OpenStudy (vickyisthesmartone123):

Blake's account started with y = 350*x + 480 which at x = 0 (before the first deposit) is y = 350*0 + 480 = 0 + 480 = $480. The equation of the line determine by the points given for Eddie's account is this: y = 475*x + 425. At x = 0, y = 475*0 + 425 = $425.

OpenStudy (vickyisthesmartone123):

Blake $480. Eddie $425.

OpenStudy (anonymous):

So $0 is what they started with ?

OpenStudy (rocketisawesome):

Sorry about that I got disconnected

OpenStudy (vickyisthesmartone123):

Hold on so what don't you get?

OpenStudy (anonymous):

Like I cant seem to figure out what they started with , I get how to solve and everything I think im just confusing myself

OpenStudy (vickyisthesmartone123):

There should be no problem because those are the answers Mrs.

OpenStudy (anonymous):

480 & 425 ?

OpenStudy (vickyisthesmartone123):

Yea i think so too...dont make it tuff on yourself kid

OpenStudy (anonymous):

Okay Ty

OpenStudy (vickyisthesmartone123):

Your welcome!:) If you need more help just tag me ill be there in a flash.

OpenStudy (chantysquirrel1129**):

Did you mention me, vicky? c:

OpenStudy (chantysquirrel1129**):

If you have something to say, say it (:

OpenStudy (chantysquirrel1129**):

http://prntscr.com/9m3jly Find the question you were going to mention me in and say what you want to say c:

OpenStudy (chantysquirrel1129**):

I don't know which question you were going to mention me in. Only you know ._.

OpenStudy (chantysquirrel1129**):

Of course I remember.

OpenStudy (chantysquirrel1129**):

Now let's get off this post, this person probably doesn't want us on their post.

OpenStudy (chantysquirrel1129**):

@I_luv_Sparkles lol this one cx

OpenStudy (chantysquirrel1129**):

I said don't spam on two post LOL I guess two is a big number for you c:

OpenStudy (anonymous):

http://prntscr.com/9m3n06

OpenStudy (chantysquirrel1129**):

LOL

OpenStudy (anonymous):

looking for something to help on

OpenStudy (anonymous):

*cough* bs *cough* lies

OpenStudy (chantysquirrel1129**):

practice what you preach, vicky c;

OpenStudy (chantysquirrel1129**):

LOL ANI XD

OpenStudy (anonymous):

*cough* liars *cough* go to *cough* hell

OpenStudy (anonymous):

i like being a feather ;3

OpenStudy (anonymous):

rite ;o

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nOPE.

OpenStudy (anonymous):

C;'

OpenStudy (anonymous):

THEN I FREAKING WON THIS...GL CHANTY

OpenStudy (anonymous):

NO I CAN'T TAKE YOUR BRIBES

OpenStudy (chantysquirrel1129**):

mods can see your deleted comments, c:

OpenStudy (chantysquirrel1129**):

chup raho

OpenStudy (chantysquirrel1129**):

I'm not even talking about spamming ._.

OpenStudy (anonymous):

YEAH

OpenStudy (chantysquirrel1129**):

LOL mhm c:

OpenStudy (vickyisthesmartone123):

OK BYE!

OpenStudy (anonymous):

http://cdn.c.photoshelter.com/img-get/I0000q6lQUDiDqSc/s/900/900/044-040-1200.jpg I FOUND YOU ON GOOGLR MCDONALDS

OpenStudy (anonymous):

GOOGLE* CX

OpenStudy (chantysquirrel1129**):

LOL no mas, can't take this. I'm dying xD

OpenStudy (anonymous):

LOL TOO CHICKEN TO FACE ANA THE STRANGE...BYE BEAVER TEETH